OpenStudy (anonymous):

Use the function ƒ(x) = 2 - 5x to fill in the blanks in the table. -2 ____________ -1 ____________ 0 _____________ 1 _______________ 2 ________________

pooja195 (pooja195):

plug in each number in place of x

pooja195 (pooja195):

5x

OpenStudy (eric_d):

Substitute the given values into the function

OpenStudy (eric_d):

@stoner_life

pooja195 (pooja195):

ƒ(x) = 2 - 5(-2) ƒ(x) = 2 - 5(-1) ƒ(x) = 2 - 5(0)

OpenStudy (anonymous):

ima need more help then this!

pooja195 (pooja195):

i set it up for u lol

pooja195 (pooja195):

you just have to solve

OpenStudy (anonymous):

all it shows is a bunch of ? marks!

pooja195 (pooja195):

refresh

pooja195 (pooja195):

ƒ(x) = 2 - 5(1) ƒ(x) = 2 - 5(2)

pooja195 (pooja195):

lets start with the first one whats -5x-2?

OpenStudy (anonymous):

idek okay .

pooja195 (pooja195):

you need to try

pooja195 (pooja195):

whats 5x2??

OpenStudy (anonymous):

damn ive already tryed this stuff aint easy like you think it is!

OpenStudy (anonymous):

its 10

pooja195 (pooja195):

ok good ƒ(x) = 10

pooja195 (pooja195):

thats the answer for the first one next: 5x-1=?

pooja195 (pooja195):

5 times -1 do you know?

pooja195 (pooja195):

ok seems your stuck whats 5 times 1?

OpenStudy (eric_d):

When x=-2, f(-2)=2-5(-2) =\(12\) rite instead of 10

pooja195 (pooja195):

i missed the 2 lol